OpenStudy (anonymous):

Help Properties of exponents? Can someone help explain these two problems to me? √9 over 6^√9? and ^4√9 √9 over ^4√9^5

Nnesha (nnesha):

\[\huge\rm \frac{ \sqrt{9} }{ 6 \sqrt{9} }\] is it ur first question ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

okay so you can cancel out sqrt {9} and sqrt{9} at the denominator right ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

so \[\huge\rm \frac{ \sqrt{9} }{ 6 \sqrt{9} }=?\]

OpenStudy (anonymous):

6?

Nnesha (nnesha):

nope what should be at the numerator ??

OpenStudy (anonymous):

3

Nnesha (nnesha):

 sqrt{9] divide by sqrt{9} = what ??